The floor of a hall 252 cm long,162cm wide is paved with equql squares .Find the largest size of marble and number required ,if only whole marbles are used.
Answer by RAY100(1637)

Let's find the factors of each dimension
.
252 factors to 2^2 * 3^2 *7
.
and 162 factors to 2*3^4
.
GCF = 2* 3^2 =18
.
252/18 = 14 pieces......162/18=9 pieces
.
total squares required is 14 *9 =126
.
check, 18*18=324,,,252*162 = 40824,,,,,40824/324 = 126,,ok